• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name r gilman

    Log entry time 12:27:21 on August31,2006

    Entry number 184101

    keyword=day shift summary



    Beam restored about 8:16 AM after Wien angle change and Mott measurement. Restart program of radiator in on hydrogen to finish up statistics.

    th_l=103.4 deg, p_l=496.1 MeV/c, th_r=24.0 deg, p_l=346.0 MeV/c
    Raster is about 4 mm high by 2.5 mm wide.

    Runs I(uA) tgt rad Notes
    3652/21914 12.5 LH2 5% 1M/7.6M
    3653/21915 12.5 LH2 5% 1M/7.4M
    3654/21916 12.5 LH2 5% 1M/7.5M

    Joe Grames in elog 1336239 reports Mott measurements of
    P = 80.6 - 85.1 %.

    Controlled access to hall to open 3/4" carbon doors.
    See ~1 mr/hr near beam calorimeter.
    On exit, do 0.1 uA beam calorimeter calibration.
    runs 3655/21917 are beam calorimeter runs.

    Changing angle requires from about 10:30 to 11:20 to move
    from 104 to 14 degrees.
    run 21918: 10 uA right arm sieve slit data, while HRSL continues to move.

    runs 3656/21918: 10 uA on LD2, raster on
    runs 3657/21919: 10 uA on LD2, raster on

    Note a few minutes before end of beam that radiator temperatre is reading high. Cause unknown, see separate halogs.

    start a cosmic run, 3658, with ps3=1, for FPP